If ship is anchored at origin of 3 dim. coordinate sys. and 2 probes are sent below the surface, 1st probe stops at (-8, -4, 6) and 2nd probe stops at
(4, -9, -2) which probe is farather from the ship?

Answer by stanbon(75887)

If ship is anchored at origin of 3 dim. coordinate sys. and 2 probes are sent below the surface, 1st probe stops at (-8, -4, 6) and 2nd probe stops at
(4, -9, -2) which probe is farther from the ship?
------------------
1st probe distance from the ship: sqrt[(-8)^2 + (-4)^2 + 6^2] = = sqrt(117)
= 10.770330...
------------
2nd probe distance from the ship: sqrt[4^2 + 9^2 + 2^2] = sqrt(101) = 10.04..
--------------------------
2nd probe is closer to the ship.
==================================
Cheers,
Stan H.